OpenStudy (anonymous):

could some one please explain how to graph the inequality −8 greater than or equal to −5x + 2 > −38?

OpenStudy (anonymous):

your goal here is to graph the values of x that satisfy the inequalty. do you know how to solve an equation?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

can you solve this 3x-5=11

OpenStudy (anonymous):

yes it Is 5.3 I think...

OpenStudy (anonymous):

use fractions

OpenStudy (anonymous):

3x/3 and 11/3???????

OpenStudy (anonymous):

? 3x-5 = 11 => 3x = 16 (adding 5 to both sides) 3x= 16 => x = 16/3 (dividing both sides by 3)

OpenStudy (anonymous):

omg oops sorry im stupid

OpenStudy (anonymous):

come on... don't you dare do that!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

so then I would divide 16 by 3?

OpenStudy (anonymous):

well, solving an inequality is very much the same with only a few differences. 1. suppose a>b, then -a<-b so if we multiply or divide by a negative number, we have to change the direction of the inequality for example... it's true that 1 < 5, but -1 > -5 

OpenStudy (anonymous):

so its 16>3? im very confused :/